OpenStudy (anonymous):

the ratio of a quarterback's completed passes to attempted passed is 6 to 7 . if he attempted 21 passes, find how many passes he completed

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

7x = 6*21 x=6*21/7 x=6*3 = 18 completed passes.

OpenStudy (phi):

completed passes to attempted passed is 6 to 7 you have a ratio of \[ \frac{6}{7}= \frac{completed}{attempted} \] he attempted 21 means replace attempted with 21 to get \[ \frac{6}{7}= \frac{completed}{21} \] "solve" for completed by multiplying both sides by 21 \[ \frac{6}{7}\cdot 21 = \frac{completed}{\cancel{21}} \cdot \cancel{21}\] cancel the 21 on the right side. simplify the left side